OpenStudy (anonymous):

∫_2^3▒(9x-2x^3 )dx evaluate

OpenStudy (anonymous):

lolwut?

OpenStudy (anonymous):

\[∫_2^3(9x-2x^3 )dx\]

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

answer = -10

OpenStudy (travis):

[9/2x^2 - 2/2*x^4] 3_0= -40.5

OpenStudy (anonymous):

um... can u explain?

OpenStudy (anonymous):

ishaan can you work this out

OpenStudy (anonymous):

you can integrate by separating integrals \[\int\limits_{2}^{3} 9x dx - \int\limits_{2}^{3}2x ^{3}dx\]

OpenStudy (anonymous):

ah yes

OpenStudy (anonymous):

ok here we go av

OpenStudy (anonymous):

\[\int_2^3 9xdx - 2x^3dx\] \[\frac{9x^2}{2} - \frac{2x^4}{4} |_2^3\]

OpenStudy (anonymous):

then: \[\int\limits_{2}^{3} 9x = 45/2\] \[\int\limits_{2}^{3}-2x ^{3}dx = -65/2\] 45/2 - 65/2 = -10

OpenStudy (travis):

oh sorry it was 3_2 i thought it was 3_0

OpenStudy (anonymous):

ah sorry there should be no = it happened by mistake

OpenStudy (anonymous):

av do you understand that

OpenStudy (anonymous):

ah typo let me write it again

OpenStudy (anonymous):

what did you mean by that?

OpenStudy (anonymous):

sorry I didnt realize that

OpenStudy (anonymous):

\[\frac{81}{2} - \frac{2 \times 81}{4} - \frac{9 \times 4}{2} + \frac{2 \times 16}{4}\]

OpenStudy (anonymous):

This will give you -18 + 8 =-10

OpenStudy (anonymous):

Diego go the Same too Medals for Diego

OpenStudy (anonymous):

Diogo. its portuguese, not spanish hehe. Thanks

OpenStudy (anonymous):

Oh sorry Diogo